OpenStudy (anonymous):

What is the difference between -5 to the 3rd power and (-5) to the 3rd power?

OpenStudy (anonymous):

no difference

OpenStudy (anonymous):

yes there is you get two different answers @satellite73

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

yes one of them you get a negative and one you get a positive I just don't remember how to do it and need help I have a test. @satellite73

OpenStudy (anonymous):

\(-5^3=-5\times 5\times 5=-125\) and \((-5)^3=(-5)\times (-5)\times (-5)=-125\) same answer either way